Trust Biomed Ltd (TBL) is a Zambian pharmaceutical company. TBL has developed, tested, and obtained final approval for a ground breaking vaccine against Covid 19. The company is now considering launching two variants of the product, an oral one or an injectable. It cannot afford the machinery to manufacture both.

The success of either variant (Variant A or B) is dependent on the actual demand for the product and how long the pandemic persists. Five (05) levels of such demand have been identified (as L1, L2, …L5).

The payoff matrix below provides a summary of the net cashflows (in USD’000 present values) expected to be received on either variant under the respective levels of demand.

L1

L2

L3

L4

L5

Variant A

2,000

2,500

3,000

3,500

4,000

Variant B

2,500

500

3,500

6,500

9,000

Prob.- Base

0.10

0.20

0.40

0.20

0.10

Prob.- Case 1

0.05

0.10

0.45

0.25

0.15

Prob.- Case 2

0.15

0.35

0.35

0.10

0.05

Required:

Evaluate the decision situation and advise TBL on which product to launch under the Base Case scenario; assuming the presence of Risk takers, Risk averse, and Rational decision makers in the management team.

Demonstrate whether this decision would remain unchanged under either the Case 1 or Case 2 scenarios.
Answer #1

Ans

The Expected Value based on probability of occurence and demand pattern is shown below for base case and case 1 and 2:

Expected Value= \sum (P(xi)*xi) where P(xi) is probability of occurence of event xi

L1 L2 L3 L4 L5 Expected Value in cashflow for Base Case Expected Value in cashflow for Case 1 Expected Value in cashflow for Case 2
Variant A 2000 2500 3000 3500 4000 3000 3175 2775
Variant B 2500 500 3500 6500 9000 3950 4725 2875
Probability Base 0.1 0.2 0.4 0.2 0.1
Probability Case 1 0.05 0.1 0.45 0.25 0.15
Probability Case 2 0.15 0.35 0.35 0.1 0.05

The excel formulas are shown below:

L1 L2 L3 L4 L5 Expected Value in cashflow for Base Case Expected Value in cashflow for Case 1 Expected Value in cashflow for Case 2
Variant A 2000 2500 3000 3500 4000 =E9*E11+F9*F11+G9*G11+H9*H11+I9*I11 =E9*E12+F9*F12+G9*G12+H9*H12+I9*I12 =E9*E13+F9*F13+G9*G13+H9*H13+I9*I13
Variant B 2500 500 3500 6500 9000 =E10*E11+F10*F11+G10*G11+H10*H11+I10*I11 =E10*E12+F10*F12+G10*G12+H10*H12+I10*I12 =E10*E13+F10*F13+G10*G13+H10*H13+I10*I13
Probability Base 0.1 0.2 0.4 0.2 0.1
Probability Case 1 0.05 0.1 0.45 0.25 0.15
Probability Case 2 0.15 0.35 0.35 0.1 0.05

Thus the variant to be launched in base case scenario is Variant B as its expected value is greater than variantA

In both Case 1 and Case2 , I advise launch of Variant B as its Expected value is greater than variant A.